# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

### v1

@@ III-B Mathematical Intuition
Statistical analysis shows that the distinction between an underfitting model, an overfitting one, and a good fit can be reflected by their bias and variance with respect to the training labels. Ideally, a model with low variance and low bias is the most desirable and usually deemed a good fit. However, during the training phase, sometimes we see a model has low variance and high bias which indicates the model predictions are not good as well as being all over the place, this is the indication of an underfitting model. It usually happens when the model is too simple to capture the complexity of the data or has not been trained for long enough to reach convergence. On the other hand, we find a model which overfits has high variance and low bias. This is because the model becomes overly complex accommodating for all the outliers in the training data and fails to generalize on unseen test data. Figure 2(a) and 2(c) show how an underfitting and overfitting model shows low and high variance respectively. Figure 2 is a visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios. The figures illustrate how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before and after the shift. In (a) the underfitted model the change is very minimal. In the (b) Good fit model the change is also relatively little, but in the (c) Overfitted model there is a large shift in the neighboring region of that data point, which reflects the over-dependence on training data.
@@
To see the impact of D2C we consider the training set $T$ divided into $k$ equal partitions, namely $\lvert T_{1}\rvert=\lvert T_{2}\rvert=...=\lvert T_{k}\rvert=\lvert T\rvert/k$. Then, we train the model $\hat{f}_{j}$ using the $j$-th data chunk, where $j\in\{1,..,k\}$. As a result, we have $k$ separately trained models. Finally, we aggregate these edge models into one central model $\hat{f}$. This process is repeated for $p$ global epochs. For the sake of simplicity, we consider each $\hat{f}_{j}$ is a regression model with weight matrix $\boldsymbol{W_{j}}=\{w_{j1},...,w_{jn}\}$ and bias $b_{j}$. For an input vector $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ the prediction of the edge model on the $t$-th epoch can be described as: $\hat{f}_{j}(\boldsymbol{x})=\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+b_{j}$. In the case of a single global epoch, we now wish to aggregate the weight matrices creating the central weight matrix $\boldsymbol{\bar{W}}=\{\bar{w}_{1},...,\bar{w}_{n}\}$. Here, the weight matrix for the $(t+1)$-th epoch is:
$\boldsymbol{\bar{W}}^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\textrm{ and }b^{t+1}=\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$ (9)
@@
$\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}$ (10)
Let $e_{jt}$ denote the empirical error of the $j$-th model while estimating some arbitrary instance on the $t$-th global epoch. We assume this error to be a random variable with normal distribution having mean zero, i.e., $e_{jt}\sim\mathcal{N}(0,s)$. Hence, $\mathop{{}\mathbb{E}}(e_{jt})=0$ and $Var(e_{jt})=s$. Let us denote the estimation of the product of two trained models using two different data chunks by $c$, i.e, $\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})=c$. So, we have:
$\mathop{{}\mathbb{E}}(e_{jt}^{2})=Var(e_{jt})+[\mathop{{}\mathbb{E}}(e_{jt})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(e_{jt}^{2})=s+0=s$ (11)
@@
$\implies Cov(\hat{f}(\boldsymbol{x})^{t}_{j},\hat{f}(\boldsymbol{x})^{t}_{l})=c-0\times 0=c$ (12)
@@
$Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=\frac{1}{k^{2}}Var\left(\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$ (13)
@@
Relevant proofs related to the derivation of Eq. 12 and 13 are provided in Appendix A. Eq.(13) provides us with an interesting insight. If the two edge models are highly correlated i.e., their predictions and empirical error are very close to each other we can assume $s\approx c$. Hence, the variance of the central model at $(t+1)$-th epoch can be reduced to: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+s(k-1))/k\approx s$. On the other hand, if the two models are very dissimilar the same expression gives us: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+0(k-1))/k\approx s/k$. As $Var(e_{jt})=Var(\hat{f}(\boldsymbol{x})^{t})=s$, each of the edge models at epoch $t$ has a variance of prediction $s$, which means if the edge models at epoch $t$ are completely different, the variation of estimate decreases by a factor of $k$ on epoch $(t+1)$. Consequently, if the edge models are identically trained the variance remains almost unchanged. This result is similar to the one proposed by Breiman et al. [25] and Bühlmann et al. [26] through the idea of ’Bagging’, which is a meta-algorithm aggregating multiple model outputs. The analysis clearly shows that our approach does not at the least worsen the problem of overfitting in this scenario but rather improves the central model on each epoch if certain conditions are met. However, in reality, the edge models are neither completely different nor exactly the same due to the relative distribution of the data subsets and their shared model architecture. So, the degree to which D2C ameliorates overfitting depends on the model architecture and the correlation among the edge models. However, it is easy to see that if the training data is contaminated by outliers and random noise, the learned parameters from the different portions of the data are more likely to be different from each other. This is because the individual outliers and noise are not highly correlated, hence distributing them among the subsets likely results in significantly different samples in the subsets, making the subsequent models different from each other. Our results show that repeating the process in every global epoch reduces overfitting in most cases. A limitation of this analysis is that deep neural networks are mostly black boxes and their architectures vary tremendously, which makes establishing a generalized mathematical representation of them extremely difficult. Hence, our assumption of a simple regression model is not an exact reflection of the actual implementation. Despite that, it describes an ideal and simple scenario where the proposed method should improve the model. Moreover, it also gave us the insight that this approach can be used in conjunction with other overfitting reduction techniques such as Dropouts(Srivastava et al., 2014) [4]. If dropouts are stacked on top of our method in every iteration of local training, the neurons are randomly removed with some probability $p$(usually $p=0.5$). This introduction of randomness ensures that edge models are different from each other increasing the chances of improvement when the models are aggregated.
IV Proposed Methodology
IV-A Creating Training Subsets
To train multiple Neural Networks parallelly using different subsets of the data, we need to divide the whole training set into multiple shards, and we also need multiple identical models that will be used to train the different subsets. These models will be denoted as Edge models in this paper. At first, the training set is randomly shuffled and then divided into multiple subsets, maintaining consistent class distribution. Each subset is then fed into a separate edge model with identical neural network architecture and trainable parameters to enable parameter averaging.
IV-B Training & Averaging of Trained Parameters
Instead of one training loop, the implementation of our method requires two loops: an inner loop for subset training which is nested within an outer loop that performs central averaging. A central model with the same architecture as the edge models is first initialized. In total, we have one Central Model and $N$ Edge models, all of which share an identical architecture, where $N$ is the number of subsets we have divided the training set into. Initially, these identical model objects are built and loaded for each training subset. Within the inner loop, each edge model is iteratively trained on its subset for a specified number of local epochs. Then, the edge models’ parameters are scaled based on their data fraction and added to a local weight list. In the outer loop, these scaled parameters are averaged, updating the central model’s parameters. Each edge model’s weights are then reinitialized to match the central model, completing a global training epoch. This process repeats over several global epochs, with weighted parameter averaging to account for the data fraction of each subset. The whole process is summarized in Figure 3, and the pseudo-code to implement D2C is provided in Algorithm 1.
Algorithm 1 Divide2Conquer

### v2

@@ 2.2. Mathematical Intuition
Statistical analysis shows that the distinction between an underfitting model, an overfitting one, and a good fit can be reflected by their bias and variance with respect to the training labels. Ideally, a model with low variance and low bias is the most desirable and usually deemed a good fit. However, during the training phase, sometimes we see a model has low variance and high bias which indicates the model predictions are not good as well as being all over the place, this is the indication of an underfitting model. It usually happens when the model is too simple to capture the complexity of the data or has not been trained for long enough to reach convergence. On the other hand, we find a model that overfits has high variance and low bias. This is because the model becomes overly complex, accommodating all the outliers in the training data, and fails to generalize to unseen test data. Figure 2(a) and 2(c) show how an underfitting and overfitting model show low and high variance, respectively. Figure 2 is a visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios. The figures illustrate how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before and after the shift. In (a) the underfitted model the change is very minimal. In the (b) Good fit model the change is also relatively little, but in the (c) Overfitted model, there is a large shift in the neighboring region of that data point, which reflects the over-dependence on training data.
@@
To see the impact of Divide2Conquer we consider the training set $T$ divided into $k$ equal partitions, namely $\lvert T_{1}\rvert=\lvert T_{2}\rvert=...=\lvert T_{k}\rvert=\lvert T\rvert/k$. Then, we train the model $\hat{f}_{j}$ using the $j$-th data chunk, where $j\in\{1,..,k\}$. As a result, we have $k$ separately trained models. Finally, we aggregate these edge models into one central model $\hat{f}$. This process is repeated for $p$ global epochs. For the sake of simplicity, we consider each $\hat{f}_{j}$ is a regression model with weight matrix $\boldsymbol{W_{j}}=\{w_{j1},...,w_{jn}\}$ and bias $b_{j}$. For an input vector $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ the prediction of the edge model on the $t$-th epoch can be described as: $\hat{f}_{j}(\boldsymbol{x})=\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+b_{j}$. In the case of a single global epoch, we now wish to aggregate the weight matrices, creating the central weight matrix $\boldsymbol{\bar{W}}=\{\bar{w}_{1},...,\bar{w}_{n}\}$. Here, the weight matrix for the $(t+1)$-th epoch is:
(9) $\boldsymbol{\bar{W}}^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\textrm{ and }b^{t+1}=\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$
@@
(10) $\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}$
Let $e_{jt}$ denote the empirical error of the $j$-th model while estimating some arbitrary instance on the $t$-th global epoch. We assume this error to be a random variable with normal distribution having mean zero, i.e., $e_{jt}\sim\mathcal{N}(0,s)$. Hence, $\mathop{{}\mathbb{E}}(e_{jt})=0$ and $Var(e_{jt})=s$. Let us denote the estimation of the product of two trained models using two different data chunks by $c$, i.e., $\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})=c$. So, we have:
(11) $\mathop{{}\mathbb{E}}(e_{jt}^{2})=Var(e_{jt})+[\mathop{{}\mathbb{E}}(e_{jt})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(e_{jt}^{2})=s+0=s$
@@
(12) $\implies Cov(\hat{f}(\boldsymbol{x})^{t}_{j},\hat{f}(\boldsymbol{x})^{t}_{l})=c-0\times 0=c$
@@
(13) $Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=\frac{1}{k^{2}}Var\left(\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$
@@
Eq.(13) provides us with an interesting insight. If the two edge models are highly correlated, i.e., their predictions and empirical error are very close to each other, we can assume $s\approx c$. Hence, the variance of the central model at $(t+1)$-th epoch can be reduced to: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+s(k-1))/k\approx s$. On the other hand, if the two models are very dissimilar the same expression gives us: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+0(k-1))/k\approx s/k$. As $Var(e_{jt})=Var(\hat{f}(\boldsymbol{x})^{t})=s$, each of the edge models at epoch $t$ has a variance of prediction $s$, which means if the edge models at epoch $t$ are completely different, the variation of estimate decreases by a factor of $k$ on epoch $(t+1)$. Consequently, if the edge models are identically trained the variance remains almost unchanged. This result is similar to the one proposed by Breiman et al. (breiman1996bagging, 2) and Bühlmann et al. (buhlmann2002analyzing, 3) through the idea of ’Bagging’, which is a meta-algorithm aggregating multiple model outputs. The analysis clearly shows that our approach does not at least worsen the problem of overfitting in this scenario, but rather improves the central model on each epoch if certain conditions are met. However, in reality, the edge models are neither completely different nor exactly the same due to the relative distribution of the data subsets and their shared model architecture. So, the degree to which D2C ameliorates overfitting depends on the model architecture and the correlation among the edge models. However, it is easy to see that if the training data is contaminated by outliers and random noise, the learned parameters from the different portions of the data are more likely to be different from each other. This is because the individual outliers and noise are not highly correlated; hence, distributing them among the subsets likely results in significantly different samples in the subsets, making the subsequent models different from each other. Our results show that repeating the process in every global epoch reduces overfitting in most cases. To provide tractable mathematical intuition for the proposed DUA-D2C/D2C framework, we have adopted a simplified linear regression model. This approach is a common and necessary simplification in theoretical analyses within the machine learning literature, utilized in seminal works such as (ghojogh2019theory, 8). While this linear model does not fully capture the non-linear complexities of the deep neural networks implemented in our experiments, it serves as a valuable and interpretable analogue. It allows us to isolate and illustrate the core statistical principles, such as variance reduction through averaging and the potential benefits of weighted aggregation, that underpin why D2C and DUA-D2C are hypothesized to improve generalization and mitigate overfitting. The insights derived from this simplified model provide a foundational understanding, which is then empirically validated through extensive experiments on complex, real-world deep learning architectures. Moreover, it also gave us the insight that this approach can be used in conjunction with other overfitting reduction techniques such as Dropouts(Srivastava et al., 2014) (srivastava2014dropout, 27). If dropouts are stacked on top of our method in every iteration of local training, the neurons are randomly removed with some probability pp(usually $p=0.5$). This introduction of randomness ensures that edge models are different from each other, increasing the chances of improvement when the models are aggregated.
2.3. Weighted Aggregation
The DUA-D2C method refines the aggregation step by introducing a dynamic, data-driven weighting scheme for the edge models, rather than a simple uniform average. Let $\alpha_{j}$ be the normalized weight assigned to the $j$-th edge model $M_{j}$ (with parameters $\boldsymbol{W_{j}}^{t},b_{j}^{t}$) at global epoch $t$. This weight $\alpha_{j}$ is derived from the model’s performance on a validation set, considering its accuracy $a_{j}$ and its inverse prediction uncertainty $u_{j}$ (scaled confidence), such that $\sum_{j=1}^{k}\alpha_{j}=1$ and $\alpha_{j}\geq 0$.
The aggregation for the central model’s parameters in DUA-D2C at epoch $(t+1)$ becomes:
(14) $\boldsymbol{\bar{W}}^{t+1}=\sum_{j=1}^{k}\alpha_{j}\boldsymbol{W_{j}}^{t}\quad\text{and}\quad b^{t+1}=\sum_{j=1}^{k}\alpha_{j}{b^{t}_{j}}$
Consequently, the output of the DUA-D2C central model at epoch $(t+1)$ is:
$\hat{f}(\boldsymbol{x})^{t+1}$ $=\sum_{j=1}^{k}\alpha_{j}\left(\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+{b^{t}_{j}}\right)$
(15) $=\sum_{j=1}^{k}\alpha_{j}\hat{f}(\boldsymbol{x})^{t}_{j}$
This is now a weighted average of the predictions of the edge models. The variance of this weighted sum can be expressed as:
(16) $Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=Var\left(\sum_{j=1}^{k}\alpha_{j}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$
$=\sum_{j=1}^{k}\alpha_{j}^{2}Var\left(\hat{f}(\boldsymbol{x})^{t}_{j}\right)+\sum_{i\neq l}\alpha_{i}\alpha_{l}Cov\left(\hat{f}(\boldsymbol{x})^{t}_{i},\hat{f}(\boldsymbol{x})^{t}_{l}\right)$
$=s\sum_{j=1}^{k}\alpha_{j}^{2}+c\sum_{i\neq l}\alpha_{i}\alpha_{l}$
assuming, as before, $Var(\hat{f}(\boldsymbol{x})^{t}_{j})=s$ for all $j$, and $Cov(\hat{f}(\boldsymbol{x})^{t}_{i},\hat{f}(\boldsymbol{x})^{t}_{l})=c$ for $i\neq l$.
The key insight from Eq. (16) is that the contribution of each edge model’s variance ($s$) and covariance ($c$) to the central model’s variance is now scaled by the square of its weight ($\alpha_{j}^{2}$) or the product of weights ($\alpha_{i}\alpha_{l}$). DUA-D2C aims to assign higher $\alpha_{j}$ values to edge models that generalize better (higher accuracy on a validation set) and are more confident (lower uncertainty).
  • Prioritizing Lower-Variance Models (Implicitly): While we assumed $Var(\hat{f}(\boldsymbol{x})^{t}_{j})=s$ for simplicity, in reality, edge models that perform poorly on the validation set or are highly uncertain might themselves exhibit higher intrinsic variance or be more influenced by noise in their data subset. By down-weighting these models (assigning smaller $\alpha_{j}$), DUA-D2C effectively reduces their contribution to the overall variance of the central model. If an edge model $M_{j}$ is deemed unreliable (low $\alpha_{j}$), its $s\cdot\alpha_{j}^{2}$ term becomes small.
  • Reducing Impact of Correlated ”Bad” Models: If several edge models are similarly poor (e.g., all overfit to noise in a similar way, leading to correlated errors, high $c$), but are identified as such by the validation process, their $\alpha_{i}\alpha_{l}$ terms will be small, thus mitigating the negative impact of correlated, poorly generalizing models.
  • Optimal Weighting for Variance Reduction: From a portfolio theory perspective in statistics, for a given set of expected returns (analogous to model quality) and a covariance matrix, there exists an optimal set of weights that minimizes variance for a target return. While DUA-D2C’s heuristic weighting (based on validation accuracy and uncertainty) is not explicitly solving such an optimization problem in each step, it intuitively steers the aggregation towards a combination that should exhibit lower variance than a simple average, especially if the quality of edge models varies significantly. If all $\alpha_{j}=1/k$, Eq. (16) reduces to Eq. (13). However, if DUA-D2C can assign larger weights to genuinely better, less noisy models, the term $\sum\alpha_{j}^{2}$ can be smaller than $1/k$ (since $\sum\alpha_{j}=1$ and $\alpha_{j}\geq 0$, $\sum\alpha_{j}^{2}\leq(\sum\alpha_{j})^{2}=1$; for uniform weights, $\sum(1/k)^{2}=k/k^{2}=1/k$). This suggests a potential for further variance reduction compared to simple D2C if the weighting scheme is effective.
Essentially, DUA-D2C attempts to perform a ”smarter” averaging. Instead of treating all edge models equally, it gives preference to those that demonstrate better generalization capabilities on data not seen during their local training. This targeted aggregation is hypothesized to lead to a central model that not only benefits from the diversity of the D2C approach but also more robustly distills the ”signal” from the ”noise” by selectively emphasizing more reliable constituent models. The goal is to achieve a central model $\hat{f}(\boldsymbol{x})^{t+1}$ that is closer to the true underlying function $f_{0}$ by minimizing not just variance but also potentially bias, if ”better” edge models are indeed less biased.
In summary, the D2C framework, by averaging parameters from models trained on data shards, offers a mechanism to reduce the variance of the final aggregated model, particularly when edge models exhibit diversity due to the distribution of noise and outliers. The DUA-D2C extension further refines this by employing a dynamic, uncertainty-aware weighting scheme. This intelligent aggregation aims to preferentially incorporate knowledge from more reliable and generalizable edge models, potentially leading to even greater variance reduction and a central model that is more robust to overfitting. While this analysis relies on simplified assumptions for complex neural networks, it provides a foundational mathematical intuition for why these ”divide and conquer” strategies, especially with adaptive aggregation, are promising for improving model generalization.
3. Proposed Methodology
The proposed Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C) framework enhances traditional distributed training paradigms by introducing an intelligent aggregation mechanism. The methodology can be broken down into distinct phases: data partitioning, edge model training, edge model evaluation and weighting, and finally, dynamic global aggregation.
3.1. Creating Training Subsets
The initial phase involves preparing the data for distributed processing. The entire training dataset $D$ is first randomly shuffled to ensure an unbiased distribution of samples. Subsequently, $D$ is partitioned into $N$ distinct, non-overlapping subsets or shards: $\{D_{1},D_{2},\dots,D_{N}\}$. A critical consideration during this partitioning is the maintenance of class balance; efforts are made to ensure that the class proportions within each subset $D_{i}$ are representative of the overall dataset $D$. Each subset $D_{i}$ is then assigned to a dedicated edge model, $M_{i}$. All $N$ edge models, along with a central model $M_{c}$, share an identical neural network architecture and are initialized with the same trainable parameters to facilitate coherent parameter aggregation.
3.2. Local Training of Edge Models
Once the data subsets are prepared and edge models initialized (typically with parameters from the central model of the previous global epoch, or randomly at the start), the local training phase commences. This phase constitutes the inner loop of the DUA-D2C process. Each edge model $M_{i}$ is trained exclusively on its assigned data subset $D_{i}$ for a predetermined number of local epochs ($E$). Standard training procedures, such as mini-batch stochastic gradient descent with a given batch size $B$ and learning rate $lr$, are employed during these local epochs. Upon completion of $E$ local epochs, each edge model $M_{i}$ possesses an updated set of parameters, $\theta^{\prime}_{i}$.
3.3. Edge Model Evaluation and Dynamic Weighting
This phase is central to the DUA-D2C approach and distinguishes it from simpler aggregation schemes. After local training, each edge model $M_{i}$ (with its updated parameters $\theta^{\prime}_{i}$) is evaluated on a common, independent validation dataset, $D_{val}$. This evaluation serves two purposes:
  1. (1)
    To assess the generalization performance of $M_{i}$ via its validation accuracy, $a_{i}$.
  2. (2)
    To quantify the predictive uncertainty of $M_{i}$, typically measured by its average prediction entropy, $\epsilon_{i}$, over all samples in $D_{val}$.
The raw inverse entropy, $inv\_e_{i}=1/(\epsilon_{i}+\delta_{e})$ (where $\delta_{e}$ is a small constant to prevent division by zero), is first computed. This value is then capped at a predefined maximum $C_{max}$ to prevent undue influence from overly confident models: $inv\_e_{cap_{i}}=\min(inv\_e_{i},C_{max})$. This capped inverse entropy is subsequently Min-Max scaled to a range of $[0,1]$ to represent a normalized confidence score, $u_{i}$:
(17) $u_{i}=\frac{inv\_e_{cap_{i}}-inv\_e_{min}}{C_{max}-inv\_e_{min}+\epsilon_{s}}$
where $inv\_e_{min}=1/(\log(K)+\delta_{e})$ is the theoretical minimum inverse entropy (for $K$ classes) and $\epsilon_{s}$ is a small stability constant. The value of $u_{i}$ is clamped to ensure it strictly lies within $[0,1]$. Finally, a composite raw scaling factor, $s_{i}$, for each edge model $M_{i}$ is calculated as a weighted sum of its validation accuracy and its normalized confidence score:
(18) $s_{i}=\lambda\cdot a_{i}+(1-\lambda)\cdot u_{i}$
Here, $\lambda$ is a tradeoff hyperparameter balancing the importance of accuracy versus confidence.
3.4. Uncertainty-Aware Global Aggregation and Parameter Update
The global aggregation phase constitutes the outer loop of DUA-D2C. The raw scaling factors $\{s_{1},s_{2},\dots,s_{N}\}$ computed in the previous step are first normalized to ensure their sum is unity, yielding the final aggregation weights $\alpha_{i}$ for each edge model $M_{i}$:
(19) $\alpha_{i}=\frac{s_{i}}{\sum_{j=1}^{N}s_{j}+\epsilon_{s}}$
The parameters of the central model, $\theta_{c}$, are then updated by taking a weighted average of the parameters $\theta^{\prime}_{i}$ from all $N$ edge models, using these dynamically determined weights $\alpha_{i}$:
(20) $\theta_{c}\leftarrow\sum_{i=1}^{N}\alpha_{i}\cdot\theta^{\prime}_{i}$
This weighted aggregation ensures that edge models that demonstrated superior generalization (higher accuracy) and greater reliability (higher confidence) on the validation set contribute more significantly to the updated central model. After the central model $M_{c}$ is updated with the new $\theta_{c}$, its parameters are disseminated back to initialize all edge models for the subsequent global epoch. This entire cycle of local training, evaluation, dynamic weighting, and global aggregation is repeated for a specified number of global epochs ($E_{global}$).
The detailed pseudo-code for the complete DUA-D2C method is provided in Algorithm 1.
Algorithm 1 Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C)